OpenStudy (anonymous):

Reduce this algebraic fraction. x-1/x^2-1 X x-1/x^2-y^2

OpenStudy (anonymous):

(x- 1) ------------------------ (x + 1)(x - y)(x + y) (x^2 - 1) = (x - 1)(x + 1) x^2 - y^2= (x - y)(x + y)